How many liters of a 40% alcohol solution must be mixed with a 65% solution to obtain 20 liters of a 50% solution?
Answer by checkley77(12844)

.65x+.40(20-x)=.50*20
.65x+8-.40x=10
.25x=10-8
.25x=2
x=2/.25
x=8 gallons of 65% solution is used.
20-8=12 gallons of 40% solution is used.
Proof:
.65*8+.40*12=.50*20
5.2+4.8=10
10=10